OpenStudy (anonymous):

Determine if the sequence is arithmetic or geometric, then find the common difference or the common ratio. 27, 9, 3, 1, 1/3, ...

OpenStudy (anonymous):

a. geomeric, common difference=1/3 b. geometric, common ratio=1/3 c. arithmetic, common ratio=1/3 d. arithmetic, common difference=1/3

OpenStudy (anonymous):

You are dividing by 3 every time. Can you figure it out now.

OpenStudy (anonymous):

well i know that lol but i dont know how to figure out if its a geometric or an arithmetic

OpenStudy (anonymous):

and i dont know what a common ratio or common difference when they say that

OpenStudy (anonymous):

ok, hang on

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

http://www.purplemath.com/modules/series3.htm I will tell you that it's geometric, can you try for the second part. The site above will help you understand.

OpenStudy (anonymous):

The info is in the second paragraph.

OpenStudy (anonymous):

Did you find the answer?

OpenStudy (anonymous):

okay thanks , ill look at it

OpenStudy (anonymous):

no prob, I get sidetracked all the time.

OpenStudy (anonymous):

lol i was on that earlier! i didnt read the whole thing but since you looked at it too im going to look. haha yeah :)

OpenStudy (anonymous):

ohhhh okay! now i know what they mean when its a common ratio or common difference. i got it now :)

OpenStudy (anonymous):

OK, so just to be clear the answer is: b. geometric, common ratio=1/3

OpenStudy (anonymous):

hahahah yayys! thats the one i picked. thanks

OpenStudy (anonymous):

your welcome